# Spatio-temporal methods for estimating subsurface ocean thermal response to tropical cyclones

### v3

@@ Abstract
Tropical cyclones (TCs), driven by heat exchange between the air and sea, pose a substantial risk to many communities around the world. Accurate characterization of the subsurface ocean thermal response to TC passage is crucial for accurate TC intensity forecasts and for an understanding of the role that TCs play in the global climate system. However, that characterization is complicated by the high-noise ocean environment, correlations inherent in spatio-temporal data, relative scarcity of in situ observations, and the entanglement of the TC-induced signal with seasonal signals. We present a general methodological framework that addresses these difficulties, integrating existing techniques in seasonal mean field estimation, Gaussian process modeling, and nonparametric regression into a functional ANOVA model. Importantly, we improve upon past work by properly handling seasonality, providing rigorous uncertainty quantification, and treating time as a continuous variable, rather than producing estimates that are binned in time. This functional ANOVA model is estimated using in situ subsurface temperature profiles from the Argo fleet of autonomous floats through a multi-step procedure, which (1) characterizes the upper ocean seasonal shift during the TC season; (2) models the variability in the temperature observations; (3) fits a thin plate spline using the variability estimates to account for heteroskedasticity and correlation between the observations. This spline fit reveals the ocean thermal response to TC passage. Through this framework, we obtain new scientific insights into the interaction between TCs and the ocean on a global scale, including a three-dimensional characterization of the near-surface and subsurface cooling along the TC storm track and the mixing-induced subsurface warming on the track’s right side.
statistical climatology,
keywords:
@@
and
1 Introduction
Tropical cyclones (TCs) occur in most oceanic basins, posing a substantial risk to many communities across the globe. This risk is exacerbated in a warming climate, in which heightened ocean heat content and sea surface temperatures (SSTs) contribute to stronger and longer lasting TCs, with intensified flooding from increased atmospheric moisture (Trenberth et al., 2018; Potter, DiMarco and Knap, 2019; Russell et al., 2020; Trepanier, 2020; Xu et al., 2020). Predominant theory holds that TCs are chiefly driven by air-sea interchange of thermal energy (e.g., Emanuel, 1986, 1999), with SSTs above 26.5C conducive to TC intensification (McTaggart-Cowan et al., 2015). The intensification of TCs generates several observable physical phenomena, including extraction of energy from the ocean surface layer and upward mixing of cooler subsurface water, both of which register as “cold wakes”, i.e., lower temperatures compared to the pre-TC passage baseline (Emanuel, 2001; Haney et al., 2012; Haakman et al., 2019). This cooling of surface waters yields a decrease in SST, possibly inhibiting further TC intensification or even maintenance of current intensity (Bender and Ginis, 2000; Shay, 2010). The subtle interplay between TC intensification and upper-ocean thermal energy is therefore important for accurate TC intensity forecasts (Mainelli et al., 2008).
@@ 1 Introduction
Although the methodological framework presented here, in particular the specific technical choices for local estimation of the seasonal mean and ocean variability, is, to our knowledge, novel to the characterization of TC-induced ocean temperature changes, the statistical framework has close connections to functional data analysis. In particular, for a fixed depth, our model decomposing each temperature observation into terms describing the TC-induced signal, a seasonal mean, and ocean variability may be viewed as a functional analysis of variance (functional ANOVA) model (Ramsay and Silverman, 2005), which has previously been used in applications as varied as medicine (Brumback and Rice, 1998; Maldonado et al., 2002), trajectory estimation (Faraway, 1997), and speech (Ramsay et al., 1996). We refer the reader to standard texts for a comprehensive treatment of functional data analysis and spline smoothing in the context of ANOVA decompositions (Ramsay and Silverman, 2005; Gu, 2013).
@@
The past work closest to our own is Cheng, Zhu and Sriver (2015), which uses an Argo profile pairing procedure to perform a temporally binned analysis of thermal changes in depth and cross-track distance. We build upon this pairing process, which takes advantage of the highly correlated nature of spatio-temporal data, to make the following contributions. First, we provide methodology to estimate the seasonal shift in upper ocean temperature over the timespan of the profile pairs considered, based upon prior oceanographic work (Ridgway, Dunn and Wilkin, 2002; Roemmich and Gilson, 2009). We find a nontrivial seasonal effect over the timespan of the profile pairs in our analysis, as most TCs occur during months when the subsurface water column experiences warming. Armed with this information, we adjust the temperature differences by removing the estimated seasonal shift. The previous analysis of Cheng, Zhu and Sriver (2015) does not perform this seasonal adjustment, which may impact their estimate of the TC-related signal at larger time lags, as the seasonal warming would not be distinguished from their TC-related signal.
@@
For these seasonally adjusted temperature differences, we further detail a method for estimating their self-variances and cross-covariances, based upon the locally stationary Gaussian process model of Kuusela and Stein (2018). The reasons for fitting these covariances prior to producing the final fits are twofold: ocean dynamics conspire to produce additional heteroskedastic noise at large time lags; and pairs of profiles, especially those nearby in space and time, can be highly correlated. These covariance estimates are then used in a thin plate spline smoother to account for the effect of the correlated, heteroskedastic observations and to provide pointwise confidence intervals. Through these steps, we attain the chief scientific contributions of characterizing the ocean response to TC passage in the continuous time realm, rather than using a binned analysis as in Cheng, Zhu and Sriver (2015), as well as properly accounting for the seasonal variation in ocean temperatures over the time scales of the paired profiles. Moreover, our estimation of within- and between-pair covariances, which were not accounted for in Cheng, Zhu and Sriver (2015), admits more accurate fits and enables rigorous uncertainty quantification. These statistical procedures are unified in terms of a functional ANOVA model, which is fit separately over a grid of twenty depth levels.
@@
The organization of the paper is as follows. Section 2 formally motivates, in oceanographic, meteorological, and climatological terms, the scientific problem of estimating the ocean thermal response to tropical cyclone passage. Section 3 introduces the tropical cyclone track data and the Argo profile data. Section 4 describes our methodology, where we begin by presenting the functional ANOVA data model and its implied data analysis pipeline. Data preprocessing is reviewed in Section 4.1, and the process for pairing profiles and projecting pairs onto TC tracks is detailed in Section 4.2. A local mean field model for capturing TC-unrelated seasonal effects is described in Section 4.3, followed by a Gaussian process model for estimating ocean variability in Section 4.4. In Section 4.5, we describe a version of the thin plate spline that allows for covariance reweighing. Section 5 describes the results of applying our methodology to the estimation of the global ocean thermal response to TCs. We conclude with a discussion of the results and future directions in Section 6.
@@ 2 Scientific context
Latent heat, the energy needed to evaporate liquid water and released when water vapor condenses, is the primary mechanism by which energy is transferred from the ocean to the atmosphere. Through latent heat transfer, the passage of a TC induces temperature decreases at the ocean surface and within the ocean mixed layer (Fisher, 1957; Leipper, 1966; Bender and Ginis, 2000; D’Asaro et al., 2007; Lin et al., 2005; Lloyd and Vecchi, 2011; Balaguru et al., 2012; Elsberry, Fraim and Trapnell, 1976; Price, 1980). Observational studies have examined the thermodynamic response of the ocean surface to TC passage via observations from survey ships or expendable bathythermographs (Price, 1980; Shay and Goni, 2000; Cione and Uhlhorn, 2003; D’Asaro et al., 2007; Dare and McBride, 2011; Haakman et al., 2019), especially before the Argo array of profiling floats was deployed.
@@
With more than two million profiles since the early 2000s, Argo floats provide unprecedented spatial and temporal coverage of the global ocean in the upper 2000 meters (Roemmich and Gilson, 2009; Riser et al., 2016; Jayne et al., 2017). Due to their autonomous nature, Argo floats are well-positioned to sample the ocean state in three dimensions before and after TC passage. Several previous studies have leveraged Argo profiles to describe TC-related changes in the upper ocean (Liu et al., 2007; Balaguru et al., 2015; Sun et al., 2012; Qu, Song and Maes, ; Cheng, Zhu and Sriver, 2015; Lin et al., 2017; Steffen and Bourassa, 2018; Trenberth et al., 2018), focusing on changes in ocean temperature, salinity, or both, in order to analyze changes in ocean density. The closest work to ours is Cheng, Zhu and Sriver (2015), who perform a binned analysis of temperature changes in time (compared to our continuous treatment of time) and do not account for a seasonal warming effect we discern in our analysis.
@@
Processes regulating the thermal recovery of cold ocean wakes associated with the passage of a TC depend on the parameters of the wake, e.g., depth, width, and wind stress (Haney et al., 2012). Once a TC passes, thermal recovery results in a net warming of the water column, a warming which increases with TC intensity (Mei et al., 2013). Globally, this net ocean warming is on the order of 0.5 PW (Sriver, Huber and Nusbaumer, 2008; Mei and Pasquero, 2012; Mei and Pasquero, 2013; Mei et al., 2013), equivalent to approximately $\nicefrac{{1}}{{6}}$ of the ocean heat transport out of the tropics (Wunsch, 2007). In the long-term mean, this added heat is either advected poleward by ocean currents (Emanuel, 2001; Korty, Emanuel and Scott, 2008) or eventually released into the atmosphere during subsequent cold seasons (Pasquero and Emanuel, 2008; Jansen, Ferrari and Mooring, 2010). For stronger TCs, part of the extra ocean warming occurs below the winter mixed-layer depth (Mei et al., 2013) and can influence global meridional heat transport. However, the transport of this heat may be primarily equatorward (Jansen and Ferrari, 2009), in which case the heat transport would not contribute to the redistribution of heat within the climate system from the warmer tropics to the colder poles. The methodological framework presented in this manuscript provides an opportunity to contribute to this ongoing discussion because it helps characterize the evolution of TC-related ocean temperature changes in greater detail, accuracy, and rigor than previously available.
@@ 3.2 Argo float data
(a) Global coverage of Argo floats (Argo Program, 2020).
(b) Argo temperature profiles
@@ 3.3 TC-centric coordinate system
To quantify the ocean thermal response to the passage of a TC, we perform a pairing procedure inspired by Cheng, Zhu and Sriver (2015). Intuitively, we seek pairs of profiles which “straddle” the passage of TC, i.e., one observation before and one observation during or after the TC passage. These Argo profile pairs are placed into a TC-centric coordinate system through a projection process, which rewrites the space-time coordinates of an Argo profile pair in reference to a passing tropical cyclone. Figure 2 illustrates this process using the TC track and profile pairs associated with Hurricane Maria (2017), and the full set of profile pairs in the TC-centric coordinate system is depicted in Figure 3(b). The pairing and projection procedures are formalized in Section 4.2.
(a) Representation in $(x_{\mathrm{lon}},x_{\mathrm{lat}})$ coordinates
(b) Representation in $(d,\tau)$ coordinates
@@
(a) Representation in $(x_{\mathrm{lon}},x_{\mathrm{lat}})$ coordinates,
@@
(b) Representation in $(d,\tau)$ coordinates, 16025 pairs total.
@@ 4 Statistical methodology
Before the detailed exposition, we provide a high-level explanation of our modeling approach, which will be further formalized as a functional ANOVA data model. First, we partition the Argo profile database into “TC profiles”, which report timestamps immediately before and after the passage of a TC, and “non-TC profiles” which are bounded away from TCs in space and time (Section 4.1.2). Using the TC profiles, we form in Section 4.2 pairs of Argo profiles, in which the two profiles “straddle” the passage of a TC, i.e., one profile occurs before (the “baseline” profile) and one profile occurs during or after the TC passage (the “signal” profile). Then, we decompose the observed temperature difference (at each pressure level) between the baseline and signal profiles in the following manner:
@@
To elucidate the estimation of $s(d,\tau)$, we describe each of the terms in (1) formally, in terms of a functional ANOVA model. Let $T$ denote an Argo temperature observation, evaluated at a fixed pressure level $z$. The temperature difference is the difference between two terms $T_{\text{signal{}}}-T_{\text{baseline{}}}$:
@@
where $m$ is a seasonal mean field and $a$ is a Gaussian process term describing non-TC ocean variability, both elucidated fully in Sections 4.3 and 4.4.
@@ 4.1.1 Interpolation and averaging
In the analysis that follows, we use the term Argo profile to refer to the collection of the twenty gridded temperature values and the vertically averaged temperature value. The partitioning, pairing, and projection procedures of Sections 4.1.2 and 4.2 operate on these collections of values on the profile level. The statistical methodology, developed in Sections 4.3, 4.4, and 4.5, considers each of the gridded pressure levels (and the vertically averaged temperature values) separately, i.e., the statistical analysis is performed twenty-one times, for each of $z=10,\dotsc,200$ and for the averaged temperatures.
@@ 4.1.2 TC profiles and non-TC profiles
We further partition the Argo profile database into two sets, which separately characterize the ocean state in the presence and absence of TCs. If an Argo profile is within $8^{\circ}$ (latitude and longitude) of a tropical cyclone observation location and within $12$ days before to $20$ days after that tropical cyclone observation, we denote it a tropical cyclone profile (TC profile). Any profile not in this set is denoted a non-tropical cyclone profile (non-TC profile). Though the non-TC profiles do not carry any of the TC signal $s(d,\tau)$, they play a crucial role in estimating the seasonal effect $m$ and ocean variability $a$, which characterize the ocean state and dynamics under the important assumption that there are no TC-related effects. These models are fully described in Sections 4.3 and 4.4.

### v4

@@ Abstract
Tropical cyclones (TCs), driven by heat exchange between the air and sea, pose a substantial risk to many communities around the world. Accurate characterization of the subsurface ocean thermal response to TC passage is crucial for accurate TC intensity forecasts and for an understanding of the role that TCs play in the global climate system. However, that characterization is complicated by the high-noise ocean environment, correlations inherent in spatio-temporal data, relative scarcity of in situ observations, and the entanglement of the TC-induced signal with seasonal signals. We present a general methodological framework that addresses these difficulties, integrating existing techniques in seasonal mean field estimation, Gaussian process modeling, and nonparametric regression into an ANOVA decomposition model. Importantly, we improve upon past work by properly handling seasonality, providing rigorous uncertainty quantification, and treating time as a continuous variable, rather than producing estimates that are binned in time. This ANOVA model is estimated using in situ subsurface temperature profiles from the Argo fleet of autonomous floats through a multi-step procedure, which (1) characterizes the upper ocean seasonal shift during the TC season; (2) models the variability in the temperature observations; (3) fits a thin plate spline using the variability estimates to account for heteroskedasticity and correlation between the observations. This spline fit reveals the ocean thermal response to TC passage. Through this framework, we obtain new scientific insights into the interaction between TCs and the ocean on a global scale, including a three-dimensional characterization of the near-surface and subsurface cooling along the TC storm track and the mixing-induced subsurface warming on the track’s right side.
@@
Tropical cyclones (TCs) occur in most oceanic basins, posing a substantial risk to many communities across the globe. This risk is exacerbated in a warming climate, in which heightened ocean heat content and sea surface temperatures (SSTs) contribute to stronger and longer lasting TCs, with intensified flooding from increased atmospheric moisture (Trenberth et al., 2018; Potter et al., 2019; Russell et al., 2020; Trepanier, 2020; Xu et al., 2020). Predominant theory holds that TCs are chiefly driven by air-sea interchange of thermal energy (e.g., Emanuel, 1986, 1999), with SSTs above 26.5C conducive to TC intensification (McTaggart-Cowan et al., 2015). The intensification of TCs generates several observable physical phenomena, including extraction of energy from the ocean surface layer and upward mixing of cooler subsurface water, both of which register as “cold wakes”, i.e., lower temperatures compared to the pre-TC passage baseline (Emanuel, 2001; Haney et al., 2012; Haakman et al., 2019). This cooling of surface waters yields a decrease in SST, possibly inhibiting further TC intensification or even maintenance of current intensity (Bender and Ginis, 2000; Shay, 2010). The subtle interplay between TC intensification and upper-ocean thermal energy is therefore important for accurate TC intensity forecasts (Mainelli et al., 2008).
@@
The underlying statistical framework can be framed in terms of a classical ANOVA decomposition of each temperature profile into a TC-induced signal, seasonal mean effect, and ocean variability. At first glance, this appears to be a spatio-temporal ANOVA model (e.g., Luo et al., 1998. However, one must carefully note that while the seasonal mean effect and ocean variability are measured in the geographic coordinates of longitude, latitude, and time of year, the TC-induced signal is parameterized in the TC-centric coordinate system. These separate coordinate systems, and the ANOVA decomposition bridging these two domains, will be formalized in the methodological development. This novel twist on the usual spatio-temporal ANOVA model allows us to pool TC-related ocean temperature observations from all TC regions for the entire TC season, and indeed across TC seasons, while still producing our estimate of the TC-induced signal in its natural TC-centric coordinate system.
@@
The past work closest to our own is Cheng et al. (2015), which uses an Argo profile pairing procedure to perform a temporally binned analysis of thermal changes in depth and cross-track distance. We build upon this pairing process, which takes advantage of the highly correlated nature of spatio-temporal data, to make the following contributions. First, we provide methodology to estimate the seasonal shift in upper ocean temperature over the timespan of the profile pairs considered, based upon prior oceanographic work (Ridgway et al., 2002; Roemmich and Gilson, 2009). We find a nontrivial seasonal effect over the timespan of the profile pairs in our analysis, as most TCs occur during months when the subsurface water column experiences warming. Armed with this information, we adjust the temperature differences by removing the estimated seasonal shift. The previous analysis of Cheng et al. (2015) does not perform this seasonal adjustment, which may impact their estimate of the TC-related signal at larger time lags, as the seasonal warming would not be distinguished from their TC-related signal.
@@
For these seasonally adjusted temperature differences, we further detail a method for estimating their self-variances and cross-covariances, based upon the locally stationary Gaussian process model of Kuusela and Stein (2018). The reasons for fitting these covariances prior to producing the final fits are twofold: ocean dynamics conspire to produce additional heteroskedastic noise at large time lags; and pairs of profiles, especially those nearby in space and time, can be highly correlated. These covariance estimates are then used in a thin plate spline smoother to account for the effect of the correlated, heteroskedastic observations and to provide pointwise confidence intervals. Through these steps, we attain the chief scientific contributions of characterizing the ocean response to TC passage in the continuous time realm, rather than using a binned analysis as in Cheng et al. (2015), as well as properly accounting for the seasonal variation in ocean temperatures over the time scales of the paired profiles. Moreover, our estimation of within- and between-pair covariances, which were not accounted for in Cheng et al. (2015), admits more accurate fits and enables rigorous uncertainty quantification. These statistical procedures are unified in terms of an ANOVA decomposition, which is fit separately over a grid of twenty depth levels.
@@
The organization of the paper is as follows. Section 1 formally motivates, in oceanographic, meteorological, and climatological terms, the scientific problem of estimating the ocean thermal response to tropical cyclone passage. Section 2 introduces the tropical cyclone track data and the Argo profile data. Section 3 describes our methodology, where we begin by presenting the ANOVA-type data model and its implied data analysis pipeline. Data preprocessing is reviewed in Section 3.1, and the process for pairing profiles and projecting pairs onto TC tracks is detailed in Section 3.2. A local mean field model for capturing TC-unrelated seasonal effects is described in Section 3.3, followed by a Gaussian process model for estimating ocean variability in Section 3.4. In Section 3.5, we describe a version of the thin plate spline that allows for covariance reweighing. Section 4 describes the results of applying our methodology to the estimation of the global ocean thermal response to TCs. We conclude with a discussion of the results and future directions in Section 12.
@@ 1 Scientific context
Latent heat, the energy needed to evaporate liquid water and released when water vapor condenses, is the primary mechanism by which energy is transferred from the ocean to the atmosphere. Through latent heat transfer, the passage of a TC induces temperature decreases at the ocean surface and within the ocean mixed layer (Fisher, 1957; Leipper, 1966; Bender and Ginis, 2000; D’Asaro et al., 2007; Lin et al., 2005; Lloyd and Vecchi, 2011; Balaguru et al., 2012; Elsberry et al., 1976; Price, 1980). Observational studies have examined the thermodynamic response of the ocean surface to TC passage via observations from survey ships or expendable bathythermographs (Price, 1980; Shay and Goni, 2000; Cione and Uhlhorn, 2003; D’Asaro et al., 2007; Dare and McBride, 2011; Haakman et al., 2019), especially before the Argo array of profiling floats was deployed.
@@
With more than two million profiles since the early 2000s, Argo floats provide unprecedented spatial and temporal coverage of the global ocean in the upper 2000 meters (Roemmich and Gilson, 2009; Riser et al., 2016; Jayne et al., 2017). Due to their autonomous nature, Argo floats are well-positioned to sample the ocean state in three dimensions before and after TC passage. Several previous studies have leveraged Argo profiles to describe TC-related changes in the upper ocean (Liu et al., 2007; Balaguru et al., 2015; Sun et al., 2012; Qu et al., 2014; Cheng et al., 2015; Lin et al., 2017; Steffen and Bourassa, 2018; Trenberth et al., 2018), focusing on changes in ocean temperature, salinity, or both, in order to analyze changes in ocean density. The closest work to ours is Cheng et al. (2015), who perform a binned analysis of temperature changes in time (compared to our continuous treatment of time) and do not account for a seasonal warming effect we discern in our analysis.
@@
Processes regulating the thermal recovery of cold ocean wakes associated with the passage of a TC depend on the parameters of the wake, e.g., depth, width, and wind stress (Haney et al., 2012). Once a TC passes, thermal recovery results in a net warming of the water column, a warming which increases with TC intensity (Mei et al., 2013). Globally, this net ocean warming is on the order of 0.5 PW (Sriver et al., 2008; Mei and Pasquero, 2012, 2013; Mei et al., 2013), equivalent to approximately $\nicefrac{{1}}{{6}}$ of the ocean heat transport out of the tropics (Wunsch, 2007). In the long-term mean, this added heat is either advected poleward by ocean currents (Emanuel, 2001; Korty et al., 2008) or eventually released into the atmosphere during subsequent cold seasons (Pasquero and Emanuel, 2008; Jansen et al., 2010). For stronger TCs, part of the extra ocean warming occurs below the winter mixed-layer depth (Mei et al., 2013) and can influence global meridional heat transport. However, the transport of this heat may be primarily equatorward (Jansen and Ferrari, 2009), in which case the heat transport would not contribute to the redistribution of heat within the climate system from the warmer tropics to the colder poles. The methodological framework presented in this manuscript provides an opportunity to contribute to this ongoing discussion because it helps characterize the evolution of TC-related ocean temperature changes in greater detail, accuracy, and rigor than previously available.
@@ 2.2 Argo float data
((a)) Global coverage of Argo floats (Argo Program, 2020).
((b)) Argo temperature profiles
@@ 2.3 TC-centric coordinate system
To quantify the ocean thermal response to the passage of a TC, we perform a pairing procedure inspired by Cheng et al. (2015). Intuitively, we seek pairs of profiles which “straddle” the passage of TC, i.e., one observation before and one observation during or after the TC passage. These Argo profile pairs are placed into a TC-centric coordinate system through a projection process, which rewrites the space-time coordinates of an Argo profile pair in reference to a passing tropical cyclone. Figure 2 illustrates this process using the TC track and profile pairs associated with Hurricane Maria (2017), and the full set of profile pairs in the TC-centric coordinate system is depicted in Figure 3(b). The pairing and projection procedures are formalized in Section 3.2.
((a)) Representation in $(x_{\mathrm{lon}},x_{\mathrm{lat}})$ coordinates
((b)) Representation in $(d,\tau)$ coordinates
@@
((a)) Representation in $(x_{\mathrm{lon}},x_{\mathrm{lat}})$ coordinates,
@@
((b)) Representation in $(d,\tau)$ coordinates, 16025 pairs total.
@@ 3 Statistical methodology
Before the detailed exposition, we provide a high-level explanation of our modeling approach, which will be further formalized as an ANOVA decomposition data model. First, we partition the Argo profile database into “TC profiles”, which report timestamps immediately before and after the passage of a TC, and “non-TC profiles” which are bounded away from TCs in space and time (Section 3.1.2). Using the TC profiles, we form in Section 3.2 pairs of Argo profiles, in which the two profiles “straddle” the passage of a TC, i.e., one profile occurs before (the “baseline” profile) and one profile occurs during or after the TC passage (the “signal” profile). Then, we decompose the observed temperature difference (at each pressure level) between the baseline and signal profiles in the following manner:
@@
To elucidate the estimation of $s(d,\tau)$, we describe each of the terms in (1) formally, in terms of an ANOVA decomposition. Let $T$ denote an Argo temperature observation, evaluated at a fixed pressure level $z$. The temperature difference is the difference between two terms $T_{\text{signal{}}}-T_{\text{baseline{}}}$:
@@
where $m$ is a seasonal mean field and $a$ is a Gaussian process term describing non-TC ocean variability, both elucidated fully in Sections 3.3 and 3.4.
@@ 3.1.1 Interpolation and averaging
In the analysis that follows, we use the term Argo profile to refer to the collection of the twenty gridded temperature values and the vertically averaged temperature value. The partitioning, pairing, and projection procedures of Sections 3.1.2 and 3.2 operate on these collections of values on the profile level. The statistical methodology, developed in Sections 3.3, 3.4, and 3.5, considers each of the gridded pressure levels (and the vertically averaged temperature values) separately, i.e., the statistical analysis is performed twenty-one times, for each of $z=10,\dotsc,200$ and for the averaged temperatures.
@@ 3.1.2 TC profiles and non-TC profiles
We further partition the Argo profile database into two sets, which separately characterize the ocean state in the presence and absence of TCs. If an Argo profile is within $8^{\circ}$ (latitude and longitude) of a tropical cyclone observation location and within $12$ days before to $20$ days after that tropical cyclone observation, we denote it a tropical cyclone profile (TC profile). Any profile not in this set is denoted a non-tropical cyclone profile (non-TC profile). Though the non-TC profiles do not carry any of the TC signal $s(d,\tau)$, they play a crucial role in estimating the seasonal effect $m$ and ocean variability $a$, which characterize the ocean state and dynamics under the important assumption that there are no TC-related effects. These models are fully described in Sections 3.3 and 3.4.